General Meeting Summary 8/27/20

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Arshak Asaturyan, Alexandre Camsonne, Bogdan Wojtsekhoski, Carlos Munoz-Camacho, Hamlet Mkrtchyan, Marie Boer, Paulo Medeiros, Vardan Tadevosyan, Rolf Ent, Vladimir Berdnikov, Tanja Horn, Charles Hyde, Ho San Ko, Julie Roche, Jacob Murphy, David H


FRAME (Carlos)

  • All fibers are cut
  • VM2000 reflector sheets have been sent to JLab - received at JLab, located in NPS cleanroom
  • currently have two tools are available at JLab - make additional ones to increase efficiency
  • total time required for one reflector shaping is ~30 min
  • Assembly to test various things
  • PCB boards assembly with PMTs and cables and fibers - testing fiber lengths
  • Cooling plates with mu metal "frame" at the back (crystals supported in front, but stacked in the back)
  • Probes to measure temperature
  • Heat exchanger + chillers
  • Assembly instructions are being prepared


CABLES (Carlos)

  • Orsay set up documentation for cables on Wiki
  • Discussion about rating of HV connectors (Samtec)
  • S. Wood checking with C. Cuevas on possible local tests
  • More recent concerns: some pins/connectors exposed -- solution: interlock system, turn off HV when box is opened (>40V must be interlocked)
  • Action Items:
  • Add link to DSG on Cables Wiki page
  • Check with Steve and Chris on status of HV connector - update at next NPS meeting
  • Send updated drawing of frame to Paulo


DSG UPDATE (Aaron)

  • HV cable fabrication ongoing and procurement of HV cable component underway
  • Voltage stability testing
  • Developing EPICS monitoring screens, e.g. environments
  • HV rating concerns discussion
  • one note: in actual experiment running low current


DESIGN/ENGINEERING (Paulo)

  • Ongoing work: Lift fixture, Slide frame
  • Released recs: platform, rails, etc.


HV DIVIDERS

  • Waiting for work to start at JLab (Fernando's team)


CRYSTALS (Vladimir)

  • Crytur crystals arriving every month and have been inspected - all good quality so far
  • SICCAS arrived - need to identify 60 crystals for NPS
  • Action Item: Keep documentation of crystals by ID


PAC RECAP OF NPS PROPOSALS

  • PAC report not yet available - comments below from the public closeout session
  • Positron DVCS
  • PAC encourages investigation
  • currently can only approve with C2 since not all information on beam is available at this time - needed to make allocation of days
  • further develop unique physics case for positrons that cannot be reached with additional electron data
  • Timelike Compton Scattering
  • not yet convinced that experiment is feasible - need more information
  • report should have details on the experimental aspects to focus on


NEXT MEETING: THURSDAY 10 SEPTEMBER AT 9AM ET